# Multiscale Neural PDE Surrogates for Prediction and Downscaling: Application to Ocean Currents

### v1

@@ Abstract
Accurate modeling of physical systems governed by partial differential equations is a central challenge in scientific computing. In oceanography, high-resolution current data are critical for coastal management, environmental monitoring, and maritime safety. However, available satellite products, such as Copernicus data for sea water velocity at $\sim$0.08° spatial resolution and global ocean models, often lack the spatial granularity required for detailed local analyses. In this work, we (a) introduce a supervised deep learning framework based on neural operators for solving PDEs and providing arbitrary resolution solutions, and (b) propose downscaling models with an application to Copernicus ocean current data. Additionally, our method can model surrogate PDEs and predict solutions at arbitrary resolution, regardless of the input resolution. We evaluated our model on real-world Copernicus ocean current data and synthetic Navier–Stokes simulation datasets.
Machine Learning, ICML
@@ 1 Introduction
Accurate and high-resolution marine current fields are foundational to numerous marine applications, coastal engineering design, and autonomous navigation. Datasets like Copernicus ocean analysis product (Copernicus Marine Service, 2025) provide global coverage at roughly $0.08^{\circ}\times 0.08^{\circ}$ (approximately 9 km in mid-latitudes), which are insufficient for applications requiring detailed local dynamics.
Downscaling methods, both dynamical and statistical, have been used to bridge this resolution gap. While dynamical downscaling, employing regional ocean models, is physically rigorous, it demands substantial computational resources, often requiring days of runtime on HPC clusters. Statistical approaches offer computational efficiency but struggle with the multiscale and non-linear nature of fluid dynamics (Kruyt et al., 2022).
@@
Neural operators, particularly Fourier Neural Operators (FNOs) (Li et al., 2020), Transolver (Wu et al., 2024), FactFormer (Li et al., 2023), and Latent Spectral Models (LSMs) (Wu et al., 2023), have demonstrated remarkable ability to learn resolution-agnostic operators governing PDEs. FNOs have been successfully applied to atmospheric and oceanographic forecasting (Sun et al., 2024).
Yang et al. (2023) applied FNOs for climate model downscaling and to solve partial differential equations (PDE) that are approximated with another numerical solver. In this work, we generalize this model to timeseries, solving PDEs at arbitrary resolutions at the same time. Further, we evaluate our method in a downscaling task with practical importance such as the Copernicus ocean current dataset.
@@
To the best of our knowledge, this is the first work to present a neural surrogate capable of accurate, resolution-agnostic prediction, where the output resolution does not depend on the input resolution.
2 Related Work
2.1 Neural Operators
2.2 Embedding Physics in Deep Learning
2.3 Arbitrary-Resolution Downscaling
To address this limitation, Yang et al. (2023) introduced an FNO-based zero-shot downscaling model that generalizes to arbitrary resolutions without retraining. This approach outperforms both traditional super-resolution models and conventional neural PDE solvers on Navier–Stokes simulations and ERA5 climate fields. However, it is important to note that, in the PDE setting, the model only performs downscaling on solutions generated by external numerical solvers. Our goal is to overcome this limitation by developing a neural operator that (a) directly solves PDEs and (b) generates solutions at arbitrary resolution, independent of input resolution.
@@ 3 Methodology
3.1 Data Sources
Navier-Stokes
$\nabla\cdot\mathbf{u}(x,t)$ $=0,$ (2)
$\omega(x,0)$ $=\omega_{0}(x),$ (3)
Following (Li et al., 2020), the forcing term is fixed as $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$, and the initial vorticity $\omega_{0}(x)$ is sampled from a Gaussian random field with mean zero and spectral decay: $\omega_{0}\sim\mathcal{N}(0,r^{3/2}(-\Delta+49I)^{-2.5})$.
Copernicus Data
@@ 3.2 Resolution-agnostic Neural Operator Framework for Downscaling
i
Figure 1: The figure (inspired by Yang et al. (2023) shows the overall structure of our Temporal/static downscaling model. The low-resolution input a goes through an optional preprocessing (gradient transformation for gradient based methods) then a neural network and an upsampling block. Then an embedding function $e(\textperiodcentered)$ is returned. Finally, a neural operator takes in $e(\textperiodcentered)$ and outputs a function which gets into a reconstruction block and an optional constraint layer.
@@
For the specific task of ocean current downscaling, we evaluated several model variants within this framework. We first introduced DUNO, which uses a Universal Neural Operator (UNO) in the neural operator block. UNO generalizes well across different PDE types and is more expressive than the standard FNO. the rest of the framework is similar to the DFNO (Yang et al., 2023).
@@ 3.3 Models for Direct PDE Solution Prediction at Multiple Scales
3.4 Training Strategy
Temporal Modeling (for PDE Surrogate Task):
Benchmarking Static Downscaling on Copernicus Data:
Models are evaluated on the task of static downscaling using Copernicus current marine data. Inputs consist of low-resolution velocity fields, and models predict high-resolution outputs (2×2$\times$2 ×, 4×4$\times$4 × and 8×8$\times$8 × downscaling). Regional subsets representing different oceanographic regimes are used to assess generalization.
Loss Functions and Normalization:
We employ both L1 (MAE) and L2 (MSE) losses during training and evaluation. For perceptual quality assessment, PSNR and SSIM metrics are also computed (details provided in the Appendix). Additionally, input channels are normalized independently using channel-wise normalization.
4 Results
@@ 4.1 Temporal Models for Multiscale PDE Solving
We first evaluate the performance of the model on the 2D incompressible Navier-Stokes dataset. The goal is to learn a surrogate that accurately simulates spatio-temporal dynamics across multiple spatial resolutions. Results for two DFNO variants—standard (Temp_DFNO) and residual-based (Temp_SpecDFNO)—are reported in Table 1.
These results underscore the effectiveness of our proposed architectures as resolution-agnostic PDE surrogates, capable of simulating fluid flow dynamics with downscaling performance comparable to DFNO2 and DFNO4 from Yang et al. (2023), while removing the dependency on a separate numerical solver.
Figure 2: Low-resolution input (row 1), predictions at $16\times 16$ and $32\times 32$ (rows 2 and 4), and corresponding ground truth (rows 3 and 5) for 2D Navier–Stokes.
@@
Table 1: Performance of Temporal_DFNO, Temporal_SpecDFNO, and DFNO models at different resolutions. DFNO-2 and DFNO-4 values are taken from (Yang et al., 2023); PSNR and SSIM for these models are not provided.
@@ 4.2 Downscaling Copernicus ocean current data
Figure 3: Ground truth vs. predictions of DFNO. Rows correspond to different output resolutions: $16\times 16$, $32\times 32$ and $64\times 64$. The first column shows the model predictions, the second shows the ground truth, and the third displays the difference between them.
Despite the absence of temporal information, the DFNO maintains coherent internal structures and sharp gradients better than other baselines.
@@
To evaluate the performance of our models on real-world oceanographic data, we used sea surface velocity fields provided by the Copernicus Marine Environment Monitoring Service. To simulate coarse observations, we downsampled the original fields using average pooling to generate inputs at resolutions corresponding to 2×, 4×, and 8× coarsening factors. The model was then tasked with reconstructing higher-resolution fields from the lowest-resolution inputs. Unlike the synthetic Navier–Stokes dataset, no temporal supervision was used; each sample was treated independently as a static snapshot.
Table 2: Downscaling to $32\times 32$
Model Loss MAE MSE PSNR SSIM
CNN_2x L1 0.424 0.19343 14.06813 0.10542
L2 0.424 0.18774 14.19797 0.12992
CNN_4x L1 0.3954 0.16950 14.64186 0.12584
L2 0.395 0.16639 14.72229 0.12872
DFNO L1 0.02134 0.00099 36.97443 0.96785
L2 0.01822 0.00059 39.23959 0.97699
DUNO L1 0.03501 0.00254 32.88501 0.91708
L2 0.03440 0.00217 33.57706 0.92083
MetaGradDFNO L1 0.01424 0.00045 40.37440 0.98574
L2 0.01370 0.00033 41.74495 0.98638
MultiGradDFNO L1 0.01465 0.00049 40.07315 0.98396
L2 0.01380 0.00034 41.58843 0.98690
SpecDFNO L1 0.03407 0.00242 33.09110 0.92286
L2 0.02451 0.00106 36.66621 0.95871
SpecDFNODiff L1 0.01836 0.00064 38.85981 0.97519
L2 0.01712 0.00049 40.07214 0.97818
Table 3: Downscaling to $64\times 64$
Model Loss MAE MSE PSNR SSIM
CNN_2x L1 0.4246 0.1945 15.3113 0.0992
L2 0.4244 0.189 15.437 0.1123
CNN_4x L1 0.3953 0.172 15.844 0.121
L2 0.3954 0.1663 15.989 0.1163
DFNO L1 0.04057 0.01749 33.15377 0.85038
L2 0.04151 0.01640 33.23573 0.84019
DUNO L1 0.04271 0.00320 32.18915 0.82181
L2 0.04251 0.00314 32.55057 0.82200
MetaGradDFNO L1 0.05493 0.00962 30.33814 0.73601
L2 0.05685 0.00851 30.10898 0.71987
MultiGradDFNO L1 0.05513 0.00851 30.04188 0.73671
L2 0.05672 0.00802 29.83308 0.71801
SpecDFNO L1 0.04283 0.00682 32.23081 0.82589
L2 0.03653 0.00655 34.16133 0.87433
SpecDFNODiff L1 0.03796 0.00367 33.97083 0.86536
L2 0.03879 0.00355 33.90497 0.86177
@@
5 Discussion
5.1 Why Marine Downscaling Matters
Fine-scale currents affect oil spill trajectory forecasts, larval dispersal models, and sediment transport. High-res reconstructions support ecological resilience and climate adaptation planning.
5.2 Model comparison
5.2.1 PDE solving at multiple resolutions
Our temporal DFNO models demonstrate strong predictive capability across multiple resolutions on the Navier–Stokes benchmark. These models not only match, but in some cases outperform the DFNO2 and DFNO4 models of Yang et al. (2023), which rely on external numerical solvers to generate low-resolution solutions before downscaling them. In contrast, our approach integrates solution generation and resolution-agnostic inference into a single neural operator framework. Thus, the proposed models has successfully combined the predictive power of a classic numerical solver as well as the physics-based downscaling ability of the DFNO.
5.2.2 Static Downscaling of Real-World Ocean Data
On Copernicus sea surface velocity fields, DFNO variants significantly outperform conventional CNN baselines at all downscaling levels. However, as the resolution gap increases (at 8×), performance deteriorates (Table 4) due to the lack of informative coarse-scale details.
2× downscaling
4× downscaling
8× downscaling
(16×16 → 128×128): Significant performance degradation occurs in all models. DUNO demonstrates good performance, and SpecDFNO and SpecDFNODiff are overall and perceptually better.
Analysis
Gradient-Enhanced Models
: MetaGradDFNO and MultiGradDFNO excel at moderate downscaling factors, leveraging multiscale information. However, their performance diminishes at higher ratios, where structural details become increasingly sparse.
Spectral Residual Methods
: SpecDFNO shows robust performance across all scales, particularly excelling at zero-shot 4× downscaling. The diffusion-enhanced variant (SpecDFNODiff) provides marginal improvements but with an increased computational overhead.
5.2.3 Limitations and Scaling Behavior
@@ 5.3 Future Work
Appendix A Training configuration
A.1 Optimizer and Hyperparameters.
A.2 Data Preprocessing
Appendix B Visualizations
@@
B.1 Downscaling and predicting PDE solutions
@@
B.2 Copernicus ocean current Downscaling

### v2

@@ Abstract
Accurate modeling of physical systems governed by partial differential equations is a central challenge in scientific computing. In oceanography, high-resolution current data are critical for coastal management, environmental monitoring, and maritime safety. However, widely used satellite products, such as Copernicus sea-surface velocity at $\sim 0.08^{\circ}$ resolution and global ocean models, often lack the spatial granularity required for detailed local analyses. We (a) introduce a supervised deep learning framework based on neural operators for solving PDEs and producing arbitrary-resolution solutions, and (b) propose downscaling models applied to Copernicus ocean current data. Additionally, our method serves as a surrogate PDE model that predicts solutions at arbitrary resolution, regardless of the input resolution. We evaluate on real-world Copernicus ocean current data and synthetic Navier–Stokes simulation datasets.
@@ 1 Introduction
Accurate and high-resolution marine current fields are foundational to numerous marine applications, coastal engineering design, and autonomous navigation. Datasets like Copernicus ocean analysis product Copernicus Marine Service (2025) provide global coverage at roughly $0.08^{\circ}\times 0.08^{\circ}$ (approximately 9 km in mid-latitudes), which are insufficient for applications requiring detailed local dynamics.
Downscaling methods, both dynamical and statistical, have been used to bridge this resolution gap. While dynamical downscaling, employing regional ocean models, is physically rigorous, it demands substantial computational resources, often requiring days of runtime on HPC clusters. Statistical approaches offer computational efficiency but struggle with the multiscale and non-linear nature of fluid dynamics Kruyt et al. (2022).
@@
Neural operators, particularly Fourier Neural Operators (FNOs) (Li et al., 2020), Transolver (Wu et al., 2024), FactFormer (Li et al., 2023), and Latent Spectral Models (LSMs) (Wu et al., 2023), have demonstrated remarkable ability to learn operators governing PDEs. FNOs have been successfully applied to atmospheric and oceanographic forecasting (Sun et al., 2024).
@@
We present, to our knowledge, one of the first PDE surrogate approaches that generates solutions at arbitrary output resolution independent of the input resolution
@@ 2.1 Resolution-agnostic Neural Operator Framework for Downscaling
Figure 1: The figure (inspired by Yang et al. (2023)) shows the overall structure of our Temporal/static downscaling model. The low-resolution input a goes through an optional preprocessing (gradient transformation for gradient based methods) then a neural network and an upsampling block. Then an embedding function $e(·)$ is returned. Finally, a neural operator takes in $e(·)$ and outputs a function which gets into a reconstruction block and an optional constraint layer.
@@
For the specific task of ocean current downscaling, we evaluated several model variants within this framework. We first introduced DUNO, which uses a U-shaped Neural Operator (UNO) in the neural operator block. UNO generalizes well across different PDE types and is more expressive than the standard FNO. the rest of the framework follows DFNO Yang et al. (2023).
@@ 2.2 Models for Direct PDE Solution Prediction at Multiple Scales
3 Experiments
3.1 Datasets and Training
The datasets, training configuration and description of the baselines can be found in Appendix B.
@@ 3.2 Temporal Models for Multiscale PDE Solving
We first evaluate the performance of the model on the 2D incompressible Navier-Stokes dataset. The goal is to learn a surrogate that accurately simulates spatio-temporal dynamics across multiple spatial resolutions (Appendix D). Results for two DFNO variants—standard (Temp_DFNO) and residual-based (Temp_SpecDFNO)—are reported in Table 1.
Our temporal models deliver accurate predictions across multiple spatial resolutions on the Navier–Stokes benchmark. They achieve competitive performance compared to DFNO-2/-4 Yang et al. (2023) while enabling multi-resolution prediction within a single network. Unlike DFNO-2/-4, which depend on an external numerical solver to generate coarse solutions prior to downscaling, our approach unifies solution generation and resolution-agnostic inference in a single framework. This combines the predictive strengths associated with classical solvers and the physics-aware downscaling of DFNO, without requiring precomputed low-resolution solutions.
Table 1: Performance of Temporal_DFNO, Temporal_SpecDFNO, and DFNO models at different resolutions. DFNO-2 and DFNO-4 values are taken from Yang et al. (2023); PSNR and SSIM for these models were not provided.
@@ 3.3 Downscaling Copernicus ocean current data
To evaluate the performance of our models on real-world oceanographic data, we used sea surface velocity fields provided by the Copernicus Marine Environment Monitoring Service. To simulate coarse observations, we downsampled the original fields using average pooling to generate inputs at resolutions corresponding to 2×, 4×, and 8× coarsening factors  (Appendix D). The model was then tasked with reconstructing higher-resolution fields from the lowest-resolution inputs. Unlike the synthetic Navier–Stokes dataset, no temporal supervision was used; each sample was treated independently as a static snapshot.
The proposed DFNO variants significantly outperform conventional CNN baselines at all downscaling levels. However, as the resolution gap increases (at 8×), performance deteriorates (Table 3) due to the lack of informative coarse-scale details.
2× downscaling
4× downscaling
8× downscaling
(16×16 → 128×128)(Appendix C) : Significant performance degradation occurs in all models. DUNO demonstrates good performance, and SpecDFNO and SpecDFNODiff are overall and perceptually better.
Table 2: Downscaling results at $32\times 32$ and $64\times 64$ .
$\times32$ $\times64$
Model Loss MAE MSE PSNR SSIM MAE MSE PSNR SSIM
CNN_2x L1 0.424 0.19343 14.06813 0.10542 0.4246 0.1945 15.3113 0.0992
L2 0.424 0.18774 14.19797 0.12992 0.4244 0.1890 15.4370 0.1123
CNN_4x L1 0.3954 0.16950 14.64186 0.12584 0.3953 0.1720 15.8440 0.1210
L2 0.395 0.16639 14.72229 0.12872 0.3954 0.1663 15.9890 0.1163
DFNO L1 0.02134 0.00099 36.97443 0.96785 0.04057 0.01749 33.15377 0.85038
L2 0.01822 0.00059 39.23959 0.97699 0.04151 0.01640 33.23573 0.84019
DUNO L1 0.03501 0.00254 32.88501 0.91708 0.04271 0.00320 32.18915 0.82181
L2 0.03440 0.00217 33.57706 0.92083 0.04251 0.00314 32.55057 0.82200
MetaGradDFNO L1 0.01424 0.00045 40.37440 0.98574 0.05493 0.00962 30.33814 0.73601
L2 0.01370 0.00033 41.74495 0.98638 0.05685 0.00851 30.10898 0.71987
MultiGradDFNO L1 0.01465 0.00049 40.07315 0.98396 0.05513 0.00851 30.04188 0.73671
L2 0.01380 0.00034 41.58843 0.98690 0.05672 0.00802 29.83308 0.71801
SpecDFNO L1 0.03407 0.00242 33.09110 0.92286 0.04283 0.00682 32.23081 0.82589
L2 0.02451 0.00106 36.66621 0.95871 0.03653 0.00655 34.16133 0.87433
SpecDFNODiff L1 0.01836 0.00064 38.85981 0.97519 0.03796 0.00367 33.97083 0.86536
L2 0.01712 0.00049 40.07214 0.97818 0.03879 0.00355 33.90497 0.86177
4 Conclusion
We show that neural-operator downscaling can deliver higher-resolution current maps from low-resolution inputs, with strong gains over CNN baselines for Copernicus currents and accurate multi-resolution temporal predictions without external solvers. These results position neural operators as practical, scalable tools for ocean current analysis, and motivate uncertainty-aware and theory-driven extensions to safely push beyond moderate downscaling.
Appendix A Related Work
A.1 Neural Operators
A.2 Embedding Physics in Deep Learning
A.3 Arbitrary-Resolution Downscaling
To address this limitation, Yang et al. (2023) introduced an FNO-based zero-shot downscaling model that generalizes to arbitrary resolutions without retraining. This approach outperforms both traditional super-resolution models and conventional neural PDE solvers on Navier–Stokes simulations and ERA5 climate fields. However, it is important to note that, in the PDE setting, the model only performs downscaling on solutions generated by external numerical solvers. We overcome this limitation by developing a neural operator that (a) directly solves PDEs and (b) generates solutions at arbitrary resolution, independent of input resolution.
Appendix B Training configuration
B.1 Data Sources
Navier-Stokes
$\nabla\cdot\mathbf{u}(x,t)$ $=0,$ (2)
$\omega(x,0)$ $=\omega_{0}(x),$ (3)
Following Li et al. (2020), the forcing term is fixed as $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$, and the initial vorticity $\omega_{0}(x)$ is sampled from a Gaussian random field with mean zero and spectral decay: $\omega_{0}\sim\mathcal{N}(0,r^{3/2}(-\Delta+49I)^{-2.5})$.
Copernicus Data
B.2 Data Preprocessing
B.3 Training Strategy
Temporal Modeling (for PDE Surrogate Task):
Benchmarking Static Downscaling on Copernicus Data:
Models are evaluated on the task of static downscaling using Copernicus current marine data. Inputs consist of low-resolution velocity fields, and models predict high-resolution outputs ($2\times$, $4\times$ and $8\times$ downscaling). Regional subsets representing different oceanographic regimes are used to assess generalization.
Loss Functions and Normalization:
We employ both L1 and L2 losses during training and evaluation. For perceptual quality assessment, PSNR and SSIM metrics are also computed (details provided in the Appendix). Additionally, input channels are normalized independently using channel-wise normalization.
B.4 Optimizer and Hyperparameters.
Appendix C Limitations
Gradient-Enhanced Models:
MetaGradDFNO and MultiGradDFNO excel at moderate downscaling factors, leveraging multiscale information. However, their performance diminishes at higher ratios, where structural details become increasingly sparse.
Spectral Residual Methods:
SpecDFNO shows robust performance across all scales, particularly excelling at zero-shot 4× downscaling. The diffusion-enhanced variant (SpecDFNODiff) provides marginal improvements but with an increased computational overhead.
Zero-shot 8 times downscaling
@@ Appendix C Limitations
Appendix D Visualizations
@@
D.1 Downscaling and predicting PDE solutions
@@
D.2 Copernicus ocean current Downscaling